Prove collision (of tennis ball) was elastic with a Velocity - time graph (Tip: calculate the distance)


1
Expert's answer
2021-05-20T18:22:52-0400

Kinematic equations for objects in free-fall:

"v=v_0-gt"

"y=y_0+v_0t-gt^2\/2"


Let "v_0=0,y_0=10"

Then:

"v=-9.8t"

"y=10-4.9t^2"

"t=\sqrt{10/4.9}=1.43\ s" -the ball falls on the ground


After collision equations become:

"v=9.8(t-1.43)"

"y=4.9(t-1.43)^2"




From graphs we can see that the ball falls on the ground after "1.43\ s" , and after the same time it rises to the height of "10\ m". So, the collision is elastic.